# Objective-Free Local Learning and Emergent Language Structure in Thinking Machines

### v1

@@ 8. Discussion
8.1. Where did the data come from?
Neither punctuation nor spacing are necessary elements of writing:
[…]
This property is not special to this example, but is a general quality of human language patterns. Historically, human writing existed without specialized tokens indicating boundary. This is perhaps best illustrated by alphabets, whose fundamental tokens are meaningless beyond their sound associations. And yet such texts could (and can) still be read. This is only possible because human language is inherently tokenizable. There exists a statistical distinction between word-forming and non-word-forming patterns. The later introduction of spaces, punctuation, and capitalization served to mark these boundaries, but they are not themselves fundamental to the structure of language. They are external annotations for a structure already present in the stream of phonetic or orthographic symbols.
The existence of tokenizable structure in language raises a deeper question: why should such structure exist at all? Conventional language models don’t address this distinction. All tokens—whether letters, spaces, punctuation, or stop tokens—are treated as elements of a single, undifferentiated vocabulary. These models assume the data has already been segmented, and that its structure reflects a meaningful set of distinctions. But this sidesteps the question of origin. The model is not responsible for discovering the structure of language, only reproducing it. By assuming the existence of a token vocabulary and training on massive amounts of pre-tokenized data, the model inherits the statistical regularities of language without explaining them. It becomes a powerful recognizer, but not a theory of how language—and its specific structure—emerges at all.
To study how such structure can emerge, we present our model with the 26 lowercase alphabetic characters as its only input features. This is not a simplification, but the intended use of the model. Learning in our framework proceeds by growing a memory structure dictated by local Hebbian updates constrained to follow a hierarchical DAG. Information is retained only when the correlations in the data align with what the local memory can represent. Presenting the model with one word at a time, or separating words by zero vectors rather than an actual token, allows it to build a memory that respects the data structure. If a space character were included as an actual token, it would be treated like any other feature, leading the model to learn correlations involving it. This would break locality, and the model would no longer distinguish between word-forming and non-word-forming sequences. Space would be learned as part of words, not as a boundary, and the accepted strings of the model would no longer correspond to linguistic units. The model would simply learn a different interpretation of the data—one that reflects statistical artifacts of the tokenization, not the underlying structure of language.
This is not a deficiency, but a reflection of how learning occurs in the model. Because it has no global objective, and no access to the meaning of individual symbols, the model aligns new information with structures that have already emerged from past correlations. When trained on unstructured noise—such as a uniform string of randomly chosen letters—the model still forms memory traces. It does so by capturing those accidental patterns in the noise that align with the allowed DAG growth. The result is a memory of accepted strings whose collective statistical properties closely match the morphological structure of natural language. In this sense, the model is capable of discovering natural linguistic patterns without data, because its internal memory structure mirrors the same constraints that shape human word formation (Eugenio, 2025). Including a space token misleads the model, not because it fails to recognize it, but because it has no way of knowing that such a symbol is meant to indicate a boundary, rather than a letter.
Thus it provides a scientific explanation for the universal tokenizability of human language, not just because it processes language data in a locally coordinated (and therefore biologically plausible) way, but because it can tell us why the data exists in the first place.
8.2. Scaling Up
Scaling up the model introduces unique challenges. Naively treating the retokenizer as a conventional tokenizer—such as byte-pair encoding (BPE) (Gage, 1994) or other subword segmentation (Schuster and Nakajima, 2012)—misses the point entirely. The retokenizer is not a preprocessing step, but a dynamic process. It functions as a compositional short-term memory system, one that stores and retrieves accepted strings by activating memory traces across the hierarchy. Discarding this hierarchical structure and collapsing all learned $n$-grams into a flat vocabulary would scale poorly and break the model’s inductive biases. The purpose of the retokenizer is not to produce tokens for input to another model, but to generate new feature bases through replay, enabling creative learning by aligning input structure with emergent memory.
While subword segmentation methods such as byte-pair encoding (BPE) (Gage, 1994) and unigram language modeling (Kudo, 2018) have become standard in NLP, they rely on statistical preprocessing over large corpora. A variety of recent approaches attempt to learn token boundaries from data, rather than relying on fixed preprocessing rules. These include segmentation-aware training using multiple subword candidates (Kudo, 2018), latent segment models that infer tokenization jointly with downstream tasks (Anonymous, 2024), and byte-level systems that learn token patches directly from raw input (Pagnoni et al., 2024; Schmidt et al., 2025). Probabilistic methods such as Bayesian word segmentation (Goldwater et al., 2009) and morphology induction algorithms like Morfessor (Creutz et al., 2006) also aim to discover subword units in an unsupervised way. Yet all of these approaches depend fundamentally on the assumption that the data contains language. That is, the model must learn to tokenize from patterns already present in the input. If no such structure exists—such as when trained on noise—these models fail to recover meaningful segmentations, because there is no statistical signal to exploit.
In contrast, the framework presented here does not learn token boundaries from data but instead defines them as a side-effect of emergent memory dynamics. It does this by forming accepted sequences through local Hebbian dynamics constrained by a hierarchical DAG. Tokenization is not assumed, but emerges as a side effect of correlation alignment. No loss is minimized, no boundaries are enforced, and no tokens are treated as intrinsically meaningful (e.g stop or space tokens). This allows the model to exhibit tokenizable structure even in the absence of language data—a capability absent in systems trained on the assumption that language structure is already embedded in the dataset.
Going beyond the limitations of short-term memory thus requires solutions which don’t ruin the emergence. Replay provides such a mechanism, as its minimal structure exploits the local dynamics to grow the model in a useful way. It suggests a framework where learned tokens are correlated in short-term memory structures, whose correlations are replayed in order to contribute yet more complex tokens into long-term memory. This problem is complicated by the lack of static architecture. This makes “scaling up” a game of assuming and exploring allowed structures, including which neurons are allowed to correlate in short-term memories, and the order and manner of replay.
Compression likely plays a critical role in scaling emergent structure. In this work, we focused on the simplest and most analytically tractable setting: binding subword patterns into word embeddings. Short-term memory need not store every $n$-gram, and the few-word examples we present are representative of general behavior. We apply an exact matrix product state (MPS) compression—equivalent to a tensor train decomposition—without truncating singular values, yielding fully disentangled feature maps and one embedding token per word. The resulting long-term memory functions as an attention-like, one-to-one key-value map over these discrete embeddings.
This exact decomposition demonstrates how structured embeddings can emerge from local memory dynamics. But the same MPS protocol naturally supports approximate compression by truncating small singular values, allowing generalization across statistically similar tokens. While we reserve empirical study of approximate compression for future work, its success in many-body physics (Gourianov et al., 2022; Schollwöck, 2011; Parker et al., 2020; Stoudenmire, 2022; Gourianov et al., 2025; Gourianov, 2022), and its growing adoption in machine learning under the name tensor train (Xu et al., 2024; Su et al., 2024; Novikov et al., 2015; Hrinchuk et al., 2020), suggest that it may be essential for scaling this framework beyond the low-data regime.
8.3. Static vs Dynamic Learning
Lastly, while the incompatibility between emergent learning and modern NLP frameworks poses challenges for scaling and experimentation, it raises a more fundamental question: Are we trying to build models of cognition – capable of continual learning, inter-generational memory, and emergent language capacity – or are we refining static pattern recognizers? These are separate goals rooted in different paradigms of intelligence.
Current language models operate in a globally optimized regime: they treat learning as a statistical estimation problem, where the architecture is fixed and the learning dynamics are frozen into a high-dimensional variational ansatz. In this sense, such models are effectively thermalized – time plays no intrinsic role in learning.
Consider that “thinking” in static models means test-time compute – but in human learning, thinking is a dynamic learning process, where observed events are reprocessed. Though advances in reinforcement learning demonstrate how a pre-trained LLM can learn by reasoning to itself (Zhao et al., 2025b, a), such zero-data learning is only possible because of the pretraining, and therefore not integral to how the model learns.
This contrasts with local event-driven learners, which require thinking for the model to grow beyond its initial locally-coordinated limitations. Locally-learned events are little more than simple replayable memories stored in an uncoordinated way. Memory retrieval is therefore essential to continued learning, not just because it’s useful to do so, but because it can do little else without further coordination.
8.4. Accident of Nature
Beyond the desire to advance technology, there is also the scientific question of human cognition. Global gradient descent presupposes (1) that a global notion of correctness exists (e.g a loss function), and (2) that neurons coordinate their updates to minimize it (via teaching signals (Maoz et al., 2020)). Such global coordination is not only incompatible with the microscopic behavior of neurons, but may not be a description of any naturally occurring microscopic system. The assumption of global coordination violates a basic assumption of physics: that nature is an unconscious collection of accidents.
Attempts to justify LLMs as biologically plausible ignore this disconnect: a black-box model that passively reflects its data distribution cannot offer a generative theory of language, only an empirical fit to it. This makes drawing scientific conclusions impossible, thus leaving fundamental questions unanswered: Why is speech segmented into words? What enforces the universality of word-forming patterns? How does linguistic structure emerge in the first place, and what makes it stable across generations? Even at the low-level explored in this work, our framework addresses these questions.
@@ 8.4. Accident of Nature
We have introduced a fully local, objective-free framework for hierarchical language learning, where compositional structure emerges from unsupervised retokenization and replay. Rather than fitting to data, the model constructs an internal, interpretable hierarchy of features, enabling smooth, bounded morphological structures to arise even in the absence of training data.
This framework provides natural mechanisms for scaling up. Through smooth hierarchical composition and replay-based memory formation, the model can build long-range correlations, compress complex structures into efficient key-value memories, and recursively retokenize features across multiple scales. In principle, this enables the emergence of increasingly complex symbolic systems from simple local interactions.
However, scaling up without a global objective is not straightforward. Without loss functions to optimize, the growth of complexity depends on the evolutionary dynamics of replay, memory, and hierarchical interactions—analogous to how structure emerges in systems like Conway’s Game of Life (Rendell, 2011, 2014; Codd, 1968) or Lenia (Chan, 2019; Plantec et al., 2023; Faldor and Cully, 2024; Chan, 2020). Progress toward richer, longer-range structure may require evolutionary strategies: introducing multiple interacting sub-networks, varying timescales of memory and replay, and selecting architectures that favor the extension of correlations over time.
A crucial distinction of this approach is that the model does not fit data – it interprets data. The system is fundamentally biased toward smooth, hierarchical DAG-like representations (Solan et al., 2005; Beavers, 2010; Beavers and Harrison, [n. d.]), which govern its construction of internal structure. This intrinsic bias explains its ability to extract natural language morphology from noise, and suggests a fundamentally different paradigm of learning—one closer to human cognition, where models continually reinterpret the world rather than merely optimizing against a data distribution. This guarantees diversity in learning, where learned errors are reinterpreted in the context of intrinsic model structure, leading to language change instead of model collapse.
Ultimately, the broader goal is to build models that can develop an entire language from scratch, without preexisting data, by autonomously constructing a hierarchy of symbols, grammar, and meaning. By allowing multiple modalities of hierarchical learning—for example, connecting emergent symbol sequences with structured perceptual patterns like the finger-counting toy example—meaning can arise at the intersection of different feature sets tied together through replay and embedding. In this way, not only words but also concepts and their cross-modal relationships could emerge spontaneously, answering the question: “How did we go from a universe without language to a universe with language?”
Figure 5. Number of unique $n$-grams in a 100-name dataset (English/Spanish mix) (dat, aset) shown as backdrop. Shaded regions mark features removed by hierarchical cuts; the solid line fits a log-normal distribution. Example words from post-cut regrowth are shown.

### v2

@@ 10. Discussion
Non-local & Non-natural
Large language models are not scientific theories of language, nor are they physical models of learning. They are non-local pattern recognizers designed to minimize a global objective. Though inspired by human learning, and even sharing some surface mechanisms such as attention, their mode of coordination is wholly artificial: each parameter receives an error signal computed from the entire system. This is equivalent to imagining every neuron equipped with a walkie-talkie, receiving global instructions [^19] This is described as the teaching signal in neuroscience literature (Schneidman2020_RandomProjections). However, we stress that the fundamental problem is in the non-local processing of information, not the signaling—i.e the presence & sensitivity of a global objective. What in the brain has the degree of coordination enough to compute, using the non-local state of the brain, individual updates for every neuron? (Nothing.) on how to change its weights. No naturally occurring microscopic phenomenon exhibits such coordination.
These models do not explain language; they passively model data. This crucial fact is often ignored in the recent scientific literature (Schrimpf&all2021; Krotov2023_AstrocyteTransformer; Hosseini&all2024; Caucheteux&all2022; Rathi&all2025_TopoLM; Lillicrap2020_BackpropBrain). For example, studies have shown that the embeddings of pre-trained LLMs can be used to predict voxel-level activity in fMRI recordings (Schrimpf&all2021; Caucheteux&all2022). Yet this correlation is not evidence that the model intrinsically captures human language structure, or is a human language model. Language arises from microscopic organization in the brain, while fMRI voxels are only coarse averages of that same structure—so any predictive correlation is expected. The LLM plays a passive role in this correlation, reproducing statistical regularities already present in the data. In a universe without language data, such models would not produce language. They are no more a theory of language than a regression model that predicts planetary motion is a theory of gravity (Vafa2025_PlanetaryLLM).
Such models not only assume the existence of language, they cannot sustain it at any level (subword or otherwise). They lack the capacity for inter-generational learning—the process by which knowledge is transmitted and reinterpreted across time. When trained repeatedly on their own outputs, they undergo model collapse (Shumailov&Gal2024nature_ModelCollapse), a behavior that appears universal to any system trained by global fitting. This makes them evolutionarily sterile: they can imitate language but cannot participate in its continuation. Children, by contrast, exhibit robust inter-generational learning (Kegl&Iwata1989_LSN; Senghas&Coppola2001_SignLanguage; Senghas1995_SignLanguage; reviewCreoles2018; Mcwhorter&Good2012_Saramaccan) despite orders of magnitude less data and computation, and yet could not compete with LLMs on modern industry benchmarks.
No amount of model scaling can overcome this deficiency. It reflects a paradigmatic gulf between modern black box AI and the cognitive models of nature. Progress in understanding the latter requires starting over, and building models which reflect naturally occurring microscopic conditions, and not just looking for the model which best fits the data.
The brain isn’t a computer
Computers, at a structural & architectural level, are non-local & symbolic. By contrast, the brain is a highly uncoordinated collection of structured accidents. Local memories lack the coordination necessary to perform calculations over large arrays. Hence why the model we present here decomposes into small tensors, despite realizing global order.
Thinking as an accidental mechanism of scaling
Nature is a collection of accidents. Naturally occurring memories are simple traces of an event, such as a thumbprint in clay. They lack the coordination to do anything other than retrieve or forget. Consequently, retrieval (thinking) is essential to continued learning, because such memories can do little else without further coordination. This is not a statement about a specific microscopic mechanism, but a broad conclusion of locality. Our contribution here is to lay down the formalism for understanding how this incidental relearning occurs, and its (completely accidental) benefits: a parallel content-addressable long-term symbolic memory with non-static computational architecture.
Unlike LLM’s where “thinking” is test-time compute for improving reasoning tasks, thinking in humans is a natural phenomenon without designed purpose or utility. While “thinking” in LLM’s provides some mechanisms for reinforcement learning (Zhao&all2025_ZDR), the training is still data-dependent & non-local, and doesn’t reflect the microscopic conditions of thinking in humans—therefore it is not really “thinking” as we personally understand it. Such terminology, while a convenient analogy, is unscientific. LLM’s don’t learn by processing (and reprocessing) experiences; and not being event-driven, likely have no experience of any relatable kind.
Retrieval is therefore the natural source of continued learning. Improving a local model does not resemble optimization in global learners. There is no fine-tuning of stored memories, because each is topologically encoded and cannot be overwritten without destroying its structure. [^20] Topological states are stable: the effect needs to be large enough to close the energy gap. Without a global objective, nothing drives the system to collectively find internal representations; forcing back-propagation through local memories would destroy their plasticity and reduce the model to a static estimator.
Scaling instead occurs through competition and accumulation: as new memories are added, they compete for activation and overlap during retrieval. As thinking generates new feature bases, new types of replay become possible—the evolving architecture means more ways in which noise can be injected to catalyze retrieval. Learning and relearning therefore unfold as a dynamic, multi-scale many-body problem. It is this physics problem that is the future of continued research in naturally occurring cognitive phenomena.
@@ 11. Conclusion
We lay down the foundations for a universal description of human language as arising from naturally occurring microscopic conditions. This takes the form of a novel paradigm of AI, which realizes globally coordinated behavior as arising from highly uncoordinated learning. We find that the uncoordinated way in which coordinated behavior is learned leads to emergent multi-scale symbolic order (organized as locally-ordered trees) & an evolving computational architecture. The local memory can store an infinite number of complex memories in a highly compressed way.
Figure 10. Number of unique $n$-grams in a 100-name dataset (English/Spanish mix) shown as backdrop. Shaded regions mark features removed by hierarchical cuts; the solid line fits a log-normal distribution. Example words from post-cut regrowth are shown.